OpenStudy (anonymous):

Elena bought 320 tropical fish for a museum display. She bought 7 times as many triggerfish as parrotfish. How many of each type of fish did she buy? Which is a system of equations to model the problem if x represents the number of parrotfish Elena bought and y represents the number of triggerfish she bought? A. x + y = 7 y = 320x B. x – y = 320 y = 7x C. x + 7y = 320 y = 7x D. x + y = 320 y = 7x

OpenStudy (anonymous):

I think it is C

OpenStudy (anonymous):

me too i think C

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

Its D x + y = 320 y = 7x sub in 7x for y x + 7x = 320 8x = 320 x = 320/8 x = 40 x + y = 320 40 + y = 320 y = 320 - 40 y = 280 That leaves parrot fish bought (x) = 40 and triggerfish bought (y) = 280